OpenStudy (anonymous):

Simplify. Assume all variables represent positive real numbers. 3√-125x^6= exact answer, using radicals as needed.

jimthompson5910 (jim_thompson5910):

\[\large \sqrt[3]{-125x^6}=((-5x^2)^3)^{\frac{1}{3}}=-5x^2\] So \[\large \sqrt[3]{-125x^6}=-5x^2\]